OpenStudy (anonymous):

Jaime is playing a game that uses numbered spaces 1-100. During one of his turns, he moves forward 4 spaces, back 6 spaces, and forward 11 spaces. He ends up on space 27. What numbered space did he begin on? A. 18 B. 20 C. 36 D. 42

OpenStudy (dumbcow):

assuming the spaces are ordered the net movement = 4-6+11 = 9 so he ends up 9 spaces ahead of original spot 27 -9 = 18

OpenStudy (anonymous):

4-6 is 10

OpenStudy (anonymous):

we'll sum the spaces he moved forward 4+11=15 now subtract the spaces he moved backward 15-6=9 we want to know the present location so to find that present location- number of locations he moved 27-9=18 so his current location is 18

OpenStudy (dumbcow):

4-6 = -2 imagine you are on a number line at 4, then go to left 6, where do you end up?

OpenStudy (anonymous):

thanks guys thanks annas and dumncow

OpenStudy (anonymous):

np thats what we do....